# Sharpness-Aware Minimization with Z-Score Gradient Filtering for Neural Networks

### v2

@@ Abstract
Generalizing well in deep neural networks remains a core challenge, particularly due to their tendency to converge to sharp minima that degrade robustness. Sharpness-Aware Minimization (SAM) mitigates this by seeking flatter minima but perturbs parameters using the full gradient, which can include statistically insignificant directions. We propose ZSharp, a simple yet effective extension to SAM that applies layer-wise Z-score normalization followed by percentile-based filtering to retain only statistically significant gradient components. This selective perturbation aligns updates with curvature-sensitive directions, enhancing generalization without requiring architectural changes. ZSharp introduces only one additional hyperparameter, the percentile threshold, and remains fully compatible with existing SAM variants. Experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet using ResNet, VGG, and Vision Transformers show that ZSharp consistently outperforms SAM and its variants in test accuracy, particularly on deeper and transformer-based models. These results demonstrate that ZSharp is a principled and lightweight improvement for sharpness-aware optimization.

### v3

@@ Abstract
Sharpness-Aware Minimization (SAM) improves neural network generalization by optimizing the worst-case loss within a neighborhood of parameters, yet it perturbs parameters using the entire gradient vector, including components with low statistical significance. We introduce ZSharp, a refined sharpness-aware optimization method that incorporates layer-wise Z-score normalization followed by percentile-based filtering. This process selects only the most statistically significant gradient components—those with large standardized magnitudes—for constructing the perturbation direction. ZSharp retains the standard two-phase SAM structure of ascent and descent while modifying the ascent step to focus on sharper, curvature-relevant directions. We evaluate ZSharp on CIFAR-10, CIFAR-100, and Tiny-ImageNet using a range of models including ResNet, VGG, and Vision Transformers. Across all architectures and datasets, ZSharp consistently achieves higher test accuracy compared to SAM, ASAM, and Friendly-SAM. These results indicate that Z-score-based gradient filtering can enhance the sharpness sensitivity of the update direction, leading to improved generalization in deep neural network training.